### v2

# Edge ideals whose all matching powers are bi-Cohen-Macaulay

Abstract.
We classify all graphs $G$ satisfying the property that all matching powers $I(G)^{[k]}$ of the edge ideal $I(G)$ are bi-Cohen-Macaulay for $1\leq k\leq\nu(G)$, where $\nu(G)$ is the maximum size of a matching of $G$.
Key words and phrases:
Bi-Cohen–Macaulay ideals, Complete graphs, Paths, Generic graphs, Squarefree powers, Matching powers.
2020 Mathematics Subject Classification:
Primary 13C05, 13C14, 13C15; Secondary 05E40
Introduction
In [17], Fløystad and Vatne introduced the concept of bi-Cohen-Macaulay simplicial complex. A simplicial complex $\Delta$ is called bi-Cohen-Macaulay, if $\Delta$ and its Alexander dual $\Delta^{\vee}$ are Cohen-Macaulay. In that paper the authors associated to each simplicial complex $\Delta$, in a natural way, a complex of coherent sheaves and showed that this complex reduces to a coherent sheaf if and only if $\Delta$ is bi-Cohen-Macaulay. Such a notion has suggested the definition of bi-Cohen-Macaulay squarefree monomial ideal.
Let $S=K[x_{1},\dots,x_{n}]$ be the standard graded polynomial ring over a field $K$ and let $I\subset S$ be a squarefree monomial ideal. We say that $I$ is Cohen-Macaulay if $S/I$ is a Cohen-Macaulay ring. Recall that $I$ can be considered as the Stanley-Reisner ideal of a simplicial complex on the vertex set $[n]=\{1,\dots,n\}$. Attached to $I$ is the Alexander dual $I^{\vee}$, which is again a squarefree monomial ideal. We say that $I$ is bi-Cohen-Macaulay (bi-CM, for short) if both $I$ and $I^{\vee}$ are Cohen-Macaulay ideals. By the Eagon-Reiner criterion [19, Theorem 8.1.9] $I$ has a linear resolution if and only if $I^{\vee}$ is Cohen-Macaulay. Hence, $I$ is bi-CM if and only if it is Cohen-Macaulay with linear resolution.
Such a notion can be revisited in graph theory. More in detail, let $G$ be a finite simple graph on the vertex set $[n]=\{1,\dots,n\}$ and let $I(G)$ be the edge ideal of $G$, that is, the squarefree monomial ideal of $S$ whose generators are the monomials $x_{i}x_{j}$ of the polynomial ring $S=K[x_{1},\dots,x_{n}]$ with $\{i,j\}$ an edge of $G$. We say that $G$ is bi-CM if $I(G)$ is bi-CM. By the Eagon-Reiner criterion, previously mentioned, it follows that a bi-CM graph $G$ is connected. Indeed, if this is not the case, then there exist induced subgraphs $G_{1}$ and $G_{2}$ of $G$ such that the vertex set $V(G)$ is the disjoint union of $V(G_{1})$ and $V(G_{2})$. It follows that $I(G)=I(G_{1})+I(G_{2})$, and the ideals $I(G_{1})$ and $I(G_{2})$ are ideals in a different set of variables. Therefore, the free resolution of $S/I(G)$ is obtained as the tensor product of the resolutions of $S/I(G_{1})$ and $S/I(G_{2})$. Thus $I(G)$ has relations of degree 444, so that $I(G)$ does not have a linear resolution.
In the last years many authors have tried to classify all bi-CM graphs. The pioneer paper is [20], where the authors gave a complete classification of bi-CM graphs, up to separation, and in the case they are bipartite or chordal.
In this article we classify the graphs $G$ which satisfy the following property: all matching powers $I(G)^{[k]}$ of the edge ideal $I(G)$ of $G$ are bi-CM for all $1\leq k\leq\nu(G)$, where $\nu(G)$ is the maximum size of a matching of $G$.
This question has been inspired by many recent articles in which special classes of graphs, whose matching powers of their edge ideals are Cohen-Macaulay, have been considered. In [5], Das, Roy and Saha proved that $I(G)^{[k]}$ is Cohen-Macaulay for all $1\leq k\leq\nu(G)$, if $G$ is a Cohen-Macaulay forest, and recently in [16], Ficarra and Moradi, have proved that all matching powers of the edge ideal of a chordal graph $G$ are Cohen-Macaulay if and only if $G$ is either a complete graph or a Cohen-Macaulay forest. Furthermore, they have proved that all matching powers of the edge ideal of a Cameron-Walker graph $G$ are Cohen-Macaulay if and only if $G$ is a complete graph on 222 or 333 vertices.
In the present paper we prove that all matching powers of the edge ideal of a finite, simple graph $G$ on $n$ non-isolated vertices are bi-CM if and only if $G$ is the complete graph $K_{n}$ or the complementary graph of a path $P_{n}$ on $n$ vertices. Our main tool is the notion of vertex splittable ideal introduced in [22, Definition 2.1].
The paper is structured as follows. Section 1 contains some notions and results useful for the development of the topic. We deeply discuss the notion of matching powers of the edge ideal of a graph and its relations with the concept of squarefree powers of a squarefree monomial ideal. The main result in the section is Proposition 1.4 that states, via the notion of principal ${\bf t}$-spread Borel ideal, that all matching powers of the edge ideal of a graph $G$ are Cohen-Macaulay if $G$ is the complete graph $K_{n}$ or the complementary graph of a path $P_{n}$ on $n$ vertices, for $n\geq 4$. This result is reversed in Section 2.
Section 2 contains our main result (Theorem 2.3) that states the classification we are looking for. We prove that if $G$ is a finite, simple graph on $n\geq 4$ non-isolated vertices, then $I(G)^{[k]}$ is bi-CM for all $1\leq k\leq\nu(G)$ if and only either $G=K_{n}$ or $G\cong P_{n}^{c}$, where $P_{n}^{c}$ is the complementary graph of a path $P_{n}$ on $n$ vertices. A key result is Lemma 2.2 that is a criterion for determine if a squarefree monomial ideal is Cohen-Macaulay.
1. Auxiliary notions and results
In this section we discuss some notions and results useful for the development of the paper.
Throughout the article the graphs $G$ considered will all be finite, simple graphs, that is, they will have no double edges and no loops. Furthermore, we assume that $G$ has no isolated vertices. The vertex set of $G$ will be denoted by $V(G)$ and we will assume that $V(G)=[n]=\{1,\ldots,n\}$, unless otherwise stated. The set of edges of $G$ will be denoted by $E(G)$.
We say that $G$ is the complete graph on $n$ vertices if $E(G)=\{\{i,j\}:1\leq i<j\leq n\}$, whereas, we say that $G$ is a path on $n$ vertices if, up to a relabeling, we have $E(G)=\{\{1,2\},\{2,3\},\dots,\{n-1,n\}\}$. As usually, a complete graph on $n$ vertices will be denoted by $K_{n}$ and a path on $n$ vertices will be denoted by $P_{n}$.
Recall that the complementary graph of a graph $G$ is the graph $G^{c}$ with vertex set $V(G^{c})=V(G)$ and with edge set $E(G^{c})=E(K_{n})\setminus E(G)$.
A $k$-matching of $G$ is a subset $M$ of $E(G)$ of size $k$ such that $e\cap e^{\prime}=\emptyset$ for all $e,e^{\prime}\in M$ with $e\neq e^{\prime}$. We denote by $V(M)$ the vertex set of $M$, that is, the set $\{i\in V(G):i\in e\ \text{for}\ e\in M\}$. The matching number of $G$, denoted by $\nu(G)$, is the maximum size of a matching of $G$.
We say that a graph $H$ is a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. A subgraph $H$ of $G$ is said to be an induced subgraph if for any two vertices $i,j$ in $H$, $\{i,j\}\in E(H)$ if and only if $\{i,j\}\in E(G)$. If $A$ is a subset of $V(G)$, the induced subgraph on $A$ is the graph with vertex set $A$ and the edge set $\{\{i,j\}:\mbox{$i,j\in A$ and $\{i,j\}\in E(G)$}\}$.
[…]
Let $1\leq k\leq\nu(G)$. We denote by $I(G)^{[k]}$ the squarefree monomial ideal generated by ${\bf x}_{V(M)}$ for all $k$-matchings $M$ of $G$. We call $I(G)^{[k]}$ the matching power of $I(G)$. If $k=1$, then $I(G)^{[1]}$ is the well-known ideal, called the edge ideal of $G$ [27], and we denote it simply by $I(G)=(x_{i}x_{j}:\{i,j\}\in E(G))$.
There is a connection of such a notion with the concept of squarefree power (see, for instance, [1]) of a squarefree monomial ideal of $S$. Let $I\subset S$ be a squarefree monomial ideal and $\mathcal{G}(I)$ be its unique minimal set of monomial generators. The $k$th squarefree power of $I$, denoted by $I^{[k]}$, is the ideal generated by the squarefree monomials of $I^{k}$. Thus $u_{1}u_{2}\cdots u_{k}$, $u_{i}\in\mathcal{G}(I)$, $i\in[k]$, belongs to $\mathcal{G}(I^{[k]})$ if and only if $u_{1},u_{2},\dots,u_{k}$ is a regular sequence. Let $\nu(I)$ be the monomial grade of $I$, i.e., the maximum among the lengths of a monomial regular sequence contained in $I$. Then $I^{[k]}\neq(0)$ if and only if $k\leq\nu(I)$. Hence, the ideal $I(G)^{[k]}$ is the $k$th squarefree power of $I(G)$ and $\nu(I(G))=\nu(G)$.
See also [1, 4, 5, 6, 7, 8, 9, 10, 11, 12, 15, 16, 21, 23, 24, 25, 26] for further studies on squarefree and matching powers.
The following result is a consequence of [10, Corollary 1.3].
Lemma 1.1.
Let $G$ be a graph and let $H$ be an induced subgraph of $G$. If $I(G)^{[k]}$ has linear resolution, then $I(H)^{[k]}$ has linear resolution too.
Following [22, Definition 2.1], a monomial ideal $I\subset S$ is called vertex splittable if it can be obtained by the following recursive procedure.
  1. (i)
    If $u$ is a monomial and $I=(u)$, $I=0$ or $I=S$, then $I$ is vertex splittable.
  2. (ii)
    If there exists a variable $x_{i}$ and vertex splittable ideals $I_{1}\subset S$ and $I_{2}\subset K[x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}]$ such that $I=x_{i}I_{1}+I_{2}$, $I_{2}\subseteq I_{1}$ and $\mathcal{G}(I)$ is the disjoint union of $\mathcal{G}(x_{i}I_{1})$ and $\mathcal{G}(I_{2})$, then $I$ is vertex splittable.
In the case (ii), the decomposition $I=x_{i}I_{1}+I_{2}$ is called a vertex splitting of $I$ and $x_{i}$ is called a splitting vertex of $I$.
The following lemma is proved in [22, Theorem 3.6, Corollary 3.8] (see also [3, Proposition 3]).
Lemma 1.2.
[…]
$I(G)=x_{n}P+I(H)$
is a vertex splitting, where $P=(x_{j}:x_{j}x_{n}\in I(G))$ and $H=G\setminus\{n\}$.
Recall that for a monomial $u\in S$, the set $\textup{supp}(u)=\{i:\ x_{i}\ \textup{divides}\ u\}$ is called the support of $u$, whereas if $I$ is a monomial ideal of $S$, the set
$\textup{supp}(I)=\bigcup_{u\in\mathcal{G}(I)}\textup{supp}(u)$
is called the support of $I$.
The next result slightly generalizes the characterization of the Cohen-Macaulay vertex splittable ideals proved in [3, Theorem 2]. The proof is verbatim the same as that of [3, Theorem 2], therefore we omit it.
Theorem 1.3.
Let $I,I_{1},I_{2}\subset S$ be monomial ideals such that $I_{2}\subseteq I_{1}$, $i\notin\textup{supp}(I_{2})$, $I=x_{i}I_{1}+I_{2}$ and $\mathcal{G}(I)=\mathcal{G}(x_{i}I_{1})\cup\mathcal{G}(I_{2})$. Furthermore, we assume that $I=x_{i}I_{1}+I_{2}$ is a Betti splitting. Then, the following statements are equivalent.
  1. (a)
    $I$ is Cohen-Macaulay.
  2. (b)
    $I_{1},I_{2}$ are Cohen-Macaulay and $\textup{depth\,}S/I_{1}=\textup{depth\,}S/(I_{2},x_{i})$.
In [13], the concept of ${\bf t}$-spread strongly stable ideal was introduced. Let $d\geq 2$ and ${\bf t}=(t_{1},\dots,t_{d-1})\in\mathbb{Z}^{d-1}_{\geq 0}$. Let $u=x_{i_{1}}\cdots x_{i_{\ell}}\in S$ with $1\leq i_{1}\leq\dots\leq i_{\ell}\leq n$ and $\ell\leq d$. We say that $u$ is ${\bf t}$-spread if $i_{j+1}-i_{j}\geq t_{j}$ for all $j=1,\dots,\ell-1$.
A monomial ideal $I\subset S$ is called ${\bf t}$-spread if $\mathcal{G}(I)$ consists of ${\bf t}$-spread monomials. A ${\bf t}$-spread ideal $I\subset S$ is called ${\bf t}$-spread strongly stable if for all ${\bf t}$-spread monomials $u\in I$ and all $i<j$ such that $x_{j}$ divides $u$ and $x_{i}(u/x_{j})$ is ${\bf t}$-spread, then $x_{i}(u/x_{j})\in I$.
Let $u\in S$ be a ${\bf t}$-spread monomial. The smallest ${\bf t}$-spread ideal containing $u$ is called the principal ${\bf t}$-spread Borel ideal generated by $u$, and is denoted by $B_{\bf t}(u)$.
If $u=x_{n-t_{d-1}}x_{n-t_{d-2}-t_{d-1}}\cdots x_{n-t_{1}-\,\cdots\,-t_{d-1}}$, then $I=B_{\bf t}(u)$ is called the ${\bf t}$-spread Veronese ideal of degree $d$ in $S$, and if $t_{1}=\dots=t_{d-1}=t$ for some $t$ then $B_{\bf t}(u)$ is called the uniform $t$-spread Veronese ideal of degree $d$ in $S$.
Proposition 1.4.
Let $G\in\{K_{n},P_{n}^{c}\}$ with $n\geq 4$. Then $I(G)^{[k]}$ is bi-CM for all $1\leq k\leq\nu(G)$. Moreover, $\textup{depth\,}S/I(K_{n})=1$ and $\textup{depth\,}S/I(P_{n}^{c})=2$.
Proof.
Note that $I(K_{n})=B_{{\bf 1}}(x_{n-1}x_{n})$ and $I(P_{n}^{c})=B_{{\bf 2}}(x_{n-2}x_{n})$, where ${\bf 1}=(1)$ and ${\bf 2}=(2)$. By [2, Theorems 2.2 and 4.3], $I(G)$ is bi-CM for $G\in\{K_{n},P_{n}^{c}\}$. Let $\mathfrak{m}=(x_{1},\dots,x_{n})$. Obviously $I(K_{n})^{[k]}=\mathfrak{m}^{[2k]}$ for all $k$, and this ideal is bi-CM.
Let $G=P_{n}^{c}$. In [11, Example 1.4], the authors proved that $I(P_{n}^{c})^{[2]}=\mathfrak{m}^{[4]}$. Here we recover such a result in a simpler way. Let $x_{i}x_{j}x_{k}x_{\ell}\in\mathfrak{m}^{[4]}$ be a monomial, with $1\leq i<j<k<\ell\leq n$. Then $x_{i}x_{k},x_{j}x_{\ell}\in I(P_{n}^{c})$ because they are ${\bf 2}$-spread monomials. Hence $u\in I(P_{n}^{c})^{[2]}$ and consequently $I(P_{n}^{c})^{[2]}=\mathfrak{m}^{[4]}$, as desired. Next, by [11, Proposition 1.3], $I(P_{n}^{c})^{[k]}=\mathfrak{m}^{[2k]}$ for all $k\geq 2$, and such ideal is bi-CM. Finally, $I(G)^{[k]}$ is bi-CM for all $k$, with $G\in\{K_{n},P_{n}^{c}\}$.
The statement about the depth follows from [13, Corollary 5.3] and the Auslander-Buchsbaum formula.
2. The classification
In this section we state and prove the main result of the paper, that is, the classification of all those graphs $G$ whose matching powers $I(G)^{[k]}$ are bi-CM, for all $1\leq k\leq\nu(G)$.
First, we note that the only graphs having up to three vertices are $K_{2}$, $P_{3}$ and $K_{3}$. Only $K_{2}$ and $K_{3}$ have the property that all their matching powers are bi-CM. Indeed, $P_{3}$ is not a Cohen-Macaulay graph.
The next definition will be useful in the sequel.
Definition 2.1.
Let $I$ and $J$ be two monomial ideals of the polynomial ring $S$. The monomial ideal defined as
$I*J=(uv\ :\ u\in\mathcal{G}(I),\,v\in\mathcal{G}(J),\,\textup{supp}(u)\cap\textup{supp}(v)=\emptyset),$
is called the matching product of $I$ and $J$.
Let ${\bf 1}=(1,\dots,1)\in\mathbb{Z}^{d}_{\geq 0}$. The next lemma will be crucial for our aim.
Lemma 2.2.
[…]
Proof.
After a suitable relabeling of the variables, we may assume that
$u=x_{n-d+1}x_{n-d+2}\cdots x_{n}.$
Then, it is immediate to see that $I=B_{\bf 1}(v)$, where $v=x_{n-d}(u/x_{n-d+1})$. It follows from [2, Theorem 4.3] (or [3, Proposition 1]) that $I$ is not Cohen-Macaulay.
Theorem 2.3.
[…]
  1. (a)
    $I(G)^{[k]}$ is bi-CM for all $1\leq k\leq\nu(G)$.
  2. (b)
    Either $G=K_{n}$ or $G\cong P_{n}^{c}$.
Proof.
(b) $\Rightarrow$ (a): Follows from Proposition 1.4.
(a) $\Rightarrow$ (b): We proceed by induction on $n\geq 4$.
Let $n=4$. It is easily checked that the only bi-CM graphs on 4 non-isolated vertices are either complete graphs or isomorphic to complements of a path on four vertices. Indeed, up to isomorphism the following seven graphs are the only graphs on 4 non-isolated vertices
[…]
It is easily seen that the only bi-CM graphs among these seven graphs are $G_{3}$ and $G_{7}$. In fact, $G_{1}$ is a Cohen-Macaulay graph which is not connected (and so $I(G_{1})$ can not have linear resolution), whereas $G_{2},G_{4},G_{5}$ and $G_{6}$ are not Cohen-Macaulay. On the other hand, $G_{3}\cong P_{4}^{c}$ and $G_{7}=K_{4}$ and, by Proposition 1.4, the statement in the case $n=4$ follows.
[…]
$I(G)=x_{n}P+I(H),$
where $H=G\setminus\{n\}$ and $P=(x_{j}:x_{j}x_{n}\in I(G))$ contains $I(H)$. Using the fact that $(x_{n}P)^{[\ell]}=0$ for $\ell\geq 2$, we can note that
$I(G)^{[k]}\ =\ (x_{n}P)*I(H)^{[k-1]}+I(H)^{[k]}\ =\ x_{n}(P*I(H)^{[k-1]})+I(H)^{[k]},$ (1)
where * is the matching product previously defined.
By [14, Corollary 3.1], $PI(H)^{k-1}$ has linear resolution. Notice that $P*I(H)^{[k-1]}$ is the squarefree part of $PI(H)^{k-1}$. Hence, by [10, Lemma 1.2], $P*I(H)^{[k-1]}$ has also linear resolution. Since both the ideals $(x_{n}P)*I(H)^{[k-1]}$ and $I(H)^{[k]}$ have linear resolution, and $\mathcal{G}(I(G)^{[k]})=\mathcal{G}((x_{n}P)*I(H)^{[k-1]})\cup\mathcal{G}(I(H)^{[k]})$, by [18, Corollary 2.4] we have that (1) is a Betti splitting. Notice that $I(H)^{[k]}\subseteq P*I(H)^{[k-1]}$ and that $n\notin\textup{supp}(I(H)^{[k]})$. So, Theorem 1.3 implies that $I(H)^{[k]}$ is Cohen-Macaulay for all $1\leq k\leq\nu(G)$. By Lemma 1.1, $I(H)^{[k]}$ has linear resolution for all $1\leq k\leq\nu(H)$. Hence, $I(H)^{[k]}$ is bi-CM for all $k$. By inductive hypothesis and after a suitable relabeling of the vertices, either $H=K_{m}$ or $H\cong P_{m}^{c}$ for some $m\leq n-1$.
Since $I(G)=x_{n}P+I(H)$ is a Betti splitting and $I(G)$ is Cohen-Macaulay, Theorem 1.3 implies that
$\textup{depth\,}\frac{S}{P}\ =\ \textup{depth\,}\frac{S}{(I(G),x_{n})}\ =\ \textup{depth\,}\frac{K[x_{1},\dots,x_{n-1}]}{I(H)}.$
Consequently,
$\mu(P)\ =\ n-\textup{depth\,}\frac{K[x_{1},\dots,x_{n-1}]}{I(H)},$ (2)
where $\mu(P)$ is the cardinality of a minimal system of generators of $P$.
Now, we distinguish the two possible cases, that is, $H=K_{m}$ and $H\cong P_{m}^{c}$.
Case 1. Assume $H=K_{m}$. By Proposition 1.4 we have
$\textup{depth\,}\frac{K[x_{1},\dots,x_{n-1}]}{I(H)}\ =\ \textup{depth\,}\frac{K[x_{1},\dots,x_{m}]}{I(K_{m})}+(n-1-m)\ =\ n-m.$
From equation (2) we obtain $\mu(P)=m$. Since
$[n]\ =\ V(G)\ =\ V(H)\cup\{n\}\cup\{i\ :\ x_{i}\in P\}\ =\ [m]\cup\{n\},$
and $I(H)=I(K_{m})\subset P$, we deduce that, up to a relabeling, $P=(x_{1},\dots,x_{m-1},x_{p})$ for some $m\leq p\leq n-1$ and either $m=n-1$ or $m=n-2$. If $m=n-1$ then $P=(x_{1},\dots,x_{n-1})$ and $G$ is the complete graph, as desired.
If otherwise $m=n-2$, then we must have $p=n-1$ and $P=(x_{1},\dots,x_{n-3},x_{n-1})$. It is immediate to see that
$I(G)^{[2]}=x_{n}P*I(K_{n-2})+I(K_{n-2})^{[2]}=x_{n}(x_{1},\dots,x_{n-1})^{[3]}+(x_{1},\dots,x_{n-2})^{[4]},$
and by the argument after equation (1) this is a Betti splitting. Now, if $n=5$, then $I(G)^{[2]}=x_{5}(x_{1},\dots,x_{4})^{[3]}$ is not Cohen-Macaulay, because it is not unmixed. Otherwise, let $n\geq 6$, then $(x_{1},\dots,x_{n-2})^{[4]}\neq(0)$. By Theorem 1.3, since $I(G)^{[2]}$ is Cohen-Macaulay, we should have
$\textup{depth\,}\frac{S}{(x_{1},\dots,x_{n-1})^{[3]}}=\textup{depth\,}\frac{S}{((x_{1},\dots,x_{n-2})^{[4]},x_{n})}.$
However, by [3, Lemma 2] the first depth is equal to 3, while the second depth is equal to 4. Hence, this case does not occur.
Case 2. Assume $H\cong P_{m}^{c}$. By Proposition 1.4, we have
$\textup{depth\,}\frac{K[x_{1},\dots,x_{n-1}]}{I(H)}\ =\ \textup{depth\,}\frac{K[x_{1},\dots,x_{m}]}{I(P_{m}^{c})}+(n-1-m)\ =\ n-m+1.$
Then, by equation (2) we get that
$\mu(P)\ =\ m-1.$ (3)
Since $[n]=V(H)\cup\{i:x_{i}\in P\}\cup\{n\}$ and $V(H)=[m]$, we have $x_{m+1},\dots,x_{n-1}\in P$. Write $P=(Q,x_{m+1},\dots,x_{n-1})$, where $Q\subseteq(x_{1},\dots,x_{m})$ is a monomial prime ideal. Since $x_{m+1},\dots,x_{n-1}\notin I(H)$ and $I(H)\subseteq P$, it follows that $I(H)\subseteq Q$. Using again Proposition 1.4 we have $\textup{height}\,I(H)=m-2$, and so $\mu(Q)\geq m-2$. Consequently $\mu(P)=\mu(Q)+n-m-1\geq n-3$. Taking into account (3) we have $m\geq n-2$. Since, by definition, $P\subseteq(x_{1},\dots,x_{n-1})$, then we have $m\leq n-1$. Hence either $m=n-2$ or $m=n-1$. We distinguish the two following cases.
Case 2.1. Let $m=n-1$. Thus $H=P_{n-1}^{c}$. For $1\leq i\leq n-1$, let $P_{i}$ be the monomial prime ideal generated by the set of variables $\{x_{1},\dots,x_{n-1}\}\setminus\{x_{i}\}$. Since $P\subseteq(x_{1},\dots,x_{n-1})$ and $\mu(P)=n-2$, we see that $P=P_{i}$ for some $1\leq i\leq n-1$. If $P=P_{1}$ or $P=P_{n-1}$, then $G^{c}$ is a path, and by Proposition 1.4, $I(G)^{[k]}$ is indeed Cohen-Macaulay for all $1\leq k\leq\nu(G)$ and (b) holds in this case.
So, it is enough to show that $P$ can not be equal to $P_{i}$ for some $2\leq i\leq n-2$. Suppose that $P=P_{i}$ for some $2\leq i\leq n-2$.
We claim that
$\mathcal{G}(I(G)^{[2]})\ =\ \mathcal{G}(\mathfrak{m}^{[4]})\setminus\{x_{i-1}x_{i}x_{i+1}x_{n}\},$ (4)
where $\mathfrak{m}=(x_{1},\dots,x_{n})$. Then Lemma 2.2 shows that $I(G)^{[2]}$ is not Cohen-Macaulay, which contradicts the assumption.
Let us show that $\mathcal{G}(I(G)^{[2]})=\mathcal{G}(\mathfrak{m}^{[4]})\setminus\{x_{i-1}x_{i}x_{i+1}x_{n}\}$ if $P=P_{i}$ for some $2\leq i\leq n-2$.
Since $I(G)=I(H)+x_{n}P$, we have that
$I(G)^{[2]}$ $=\ I(H)^{[2]}+I(H)*(x_{n}P)+(x_{n}P)^{[2]}$ (5)
$=\ (x_{1},\dots,x_{n-1})^{[4]}+x_{n}(P*I(H)),$
where we have used the fact that $(x_{n}P)^{[2]}=(0)$.
By (5), all monomials of $\mathcal{G}(\mathfrak{m}^{[4]})$ which are not divided by $x_{n}$ belong to $\mathcal{G}(I(G)^{[2]})$. Let $u=x_{j}x_{k}x_{\ell}x_{n}$ be a squarefree monomial divided by $x_{n}$, with $1\leq j<k<\ell<n$. Next, we show that $u\in\mathcal{G}(I(G)^{[2]})$ if and only if $u\neq x_{i-1}x_{i}x_{i+1}x_{n}$. This will prove equation (4).
If none of the integers $j,k,\ell$ is equal to $i$, then $\{j,\ell\}\in E(G)$ because $\ell\geq j+2$ and $I(H)=I(P_{n-1}^{c})\subset I(G)$. Moreover $\{k,n\}\in E(G)$ because $x_{k}\in P$. Then, we have $u=(x_{j}x_{\ell})(x_{k}x_{n})\in I(G)^{[2]}$, as desired.
Suppose now that one of the integers $j,k,\ell$ is equal to $i$.
If $j=i$ or $\ell=i$, then in both cases $\{j,\ell\}\in E(G)$ because $\ell\geq j+2$. As before, $x_{k}\in P$ since $k\neq i$ and so $\{k,n\}\in E(G)$. Hence $u=(x_{j}x_{\ell})(x_{k}x_{n})\in I(G)^{[2]}$, once again.
Let $k=i$. Then $1\leq j\leq i-1$ and $i+1\leq\ell\leq n-1$.
Suppose that $j<i-1$. Then $\{j,i\}\in E(H)\subset E(G)$ because $i\geq j+2$. Since $\ell\neq i$, we have $x_{\ell}\in P$. Hence $\{\ell,n\}\in E(G)$, and so $u=(x_{j}x_{i})(x_{\ell}x_{n})\in I(G)^{[2]}$.
Similarly, if $\ell>i+1$, then $\{i,\ell\}\in E(H)\subset E(G)$ because $\ell\geq i+2$. Moreover $x_{j}\in P$ since $j\neq i$, and so $\{j,n\}\in E(G)$. Hence $u=(x_{i}x_{\ell})(x_{j}x_{n})\in I(G)$.
[…]
Case 2.2. Let $m=n-2$. We will show that this case can never occur, and this will conclude the proof.
Let $n$ be odd, say $n=2k+1$ for some $k\geq 2$. Since $I(G)^{[\nu(G)]}$ is Cohen-Macaulay by assumption, by [16, Theorem 1.8(b)] we have that $\nu(G)=k$ and $I(G)^{[k]}=\mathfrak{m}^{[2k]}$. In particular, $u=x_{1}\cdots x_{2k}\in\mathcal{G}(I(G)^{[k]})$. Notice that
$I(G)^{[k]}=I(P_{2k-1}^{c})^{[k]}+I(P_{2k-1}^{c})^{[k-1]}*(x_{n}P)=I(P_{2k-1}^{c})^{[k-1]}*(x_{n}P),$
because $\nu(P^{c}_{2k-1})=k-1$. Hence all minimal monomial generators of $I(G)^{[k]}$ are divided by $x_{n}$, and so $u\notin\mathcal{G}(I(G)^{[k]})$. A contradiction.
Now, let $n$ be even, say $n=2k$ with $k\geq 3$. Then $k-1\geq 2$, $\nu(G)=k$ and by the proof of Proposition 1.4 we have
$I(G)^{[k-1]}$ $=\ I(P^{c}_{2k-2})^{[k-1]}+(x_{n}P)*I(P^{c}_{2k-2})^{[k-2]}$
$=\ (x_{1}\cdots x_{2k-2})+x_{n}(P*I(P^{c}_{2k-2})^{[k-2]}).$
Since $\mu(P)=n-3$, we can find $1\leq i\leq n-2$ with $x_{i}\notin P$. We claim that $Q=(x_{i},x_{n})$ is a minimal prime ideal of $I(G)^{[k-1]}$. Indeed, from the above decomposition it is clear that $I(G)^{[k-1]}\subseteq Q$. Notice that $(x_{n})$ does not contain $I(G)^{[k-1]}$ because $x_{n}$ does not divide $x_{1}\cdots x_{2k-2}$ and $(x_{i})$ does not contain $I(G)^{[k-1]}$ because $x_{i}$ does not divide $x_{n}x_{n-1}u$ for some $u\in\mathcal{G}(I(P^{c}_{2k-2})^{[k-2]})$ with $x_{i}$ not dividing $u$. It is possible to find such a monomial $u$ because $\nu(P^{c}_{2k-2})=k-1>k-2$.
[…]
We expect that any uniform $t$-spread Veronese ideal of degree $d\geq t$ has the property that all its squarefree powers are bi-CM.
Acknowledgment. A. Ficarra was partly supported by the Grant JDC2023-051705-I funded by MICIU/AEI/10.13039/501100011033 and by the FSE+. Moreover, both the authors acknowledge support of the GNSAGA of INdAM (Italy).